find the product 16x5,28x^6
Answer by richwmiller(17219)

If I am reading the problem correctly, you want to multiply 16x5 and 28x^6

16x^5 *28x^6
16*28*x^5*x^6
think about it x^2*x^3=32
x^1=x
x^2=x*x
x^3=x*x*x
we want x^2*x^3=x*x * x*x*x= x^5
let's test that with 2
2^2=4
2^3=8
4*8=32
2^4=14
2^5=32
32=2^5

so when we multiply x's we add the exponents
so
x^5*x^6=x^11
16*28*x^11